• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 16:06:24 on November 9,2009

    Entry number 298589

    keyword=day shift summary

    11/09/09 day shift summary
    
    Shift worksers: J. P. Chen (SL), R. Michaels (TO), Nuruzzaman (3rd).
    
    8:00 production with 100 uA on LD2 with right HRS.
    Left dipole still down. 
    Run Happex 13850 7:28-8:35.
     Right 4889 7:28-8:20 5M, 4890 8:21-9:21 0.9M (no beam after 8:35)
    
    8:30 beam off for Moller quads cycle.
    9:00 moved target to EMPTY, MCC do FOTP.
    9:10-9:50 setting controlled access.
    9:50 Controlled access for
    1) Left dipole and Q1 work Jack and Techs
    2) PVDIS DAQ work: Ramesh, Kai, (later, Bob and Dingcheng)
    3) M. Friend checking Compton (trace cables).
    
    13:00 Right Q3 and Q1 tripped due to low helium level.
    15:00 Helium level recoveded. Right Q1/Q3 reset, Q3 cycled.
    
    16:00 still controlled access. People are all out. ready to go back to beam 
    permit.